# HOLE: Homological Observation of Latent Embeddings for Neural Network Interpretability

### v1

@@ 4.5.1 Pruning
Pruning is a model compression technique that helps sparsify a network by systematically removing redundant or less important parameters by setting them to zero [41, 28]. The fundamental principle behind pruning is based on the observation that many neural networks are over-parameterized [21], containing numerous weights that contribute minimally to the final output. By identifying and removing these parameters, we can achieve significant reductions in model size and computational requirements while maintaining comparable performance.
Pruning typically involves three main steps. First, during Importance Scoring, each parameter is assigned a score reflecting its contribution to the network’s performance. Next, during Selection, parameters with scores below a certain threshold are marked for removal. Finally, during Removal, the selected parameters are set to zero or completely eliminated from the network structure [7].
There are several variants of pruning strategies. With magnitude-based pruning, parameters are ranked by their absolute values, with smaller weights considered less important and removed first [27]. Structured pruning removes entire channels, filters, or layers [42, 44], while unstructured pruning removes individual weights regardless of their position in the network architecture [47]. Finally, global pruning considers all parameters across the entire network when making removal decisions [27], while layer-wise pruning makes decisions independently within each layer [52].
@@
We apply magnitude-based unstructured pruning [27] to remove weights below a threshold $\tau$, effectively setting $w_{ij}=0$ for $|w_{ij}|<\tau$. This sparsification process reduces the effective dimensionality of activation spaces and may eliminate critical topological connections. We evaluate pruning ratios from 10% to 90% and analyze how the resulting sparse networks maintain or lose persistent homological features across layers. We implement two distinct pruning strategies to analyze their differential impact on topological structure:
@@ 4.5.2 Quantization
Quantization is a compression technique that reduces the precision of numerical representations used in neural networks, typically converting floating-point weights and activations to lower bit-width representations [22]. This process significantly reduces memory requirements and computational complexity, making models more suitable for deployment on resource-constrained devices such as mobile phones and embedded systems [32].
The quantization process involves mapping continuous floating-point values to a discrete set of quantized values [37]. The most common approach is uniform quantization, where the range of possible values is divided into equally-spaced intervals. For a given floating-point value $x$, the quantized representation $q$ is computed as: $q=\text{round}\left(\frac{x-z}{s}\right)$, where $s$ is the scale factor that determines the spacing between quantized levels, and $z$ is the zero-point that ensures the quantization scheme can represent zero exactly [32].
Several quantization strategies exist. Post-training quantization is applied after the model has been fully trained, requiring no retraining but potentially leading to accuracy degradation [3, 55]. Quantization-aware training incorporates quantization effects during the training process, typically achieving better accuracy preservation [32, 19]. Dynamic quantization determines scale and zero-point parameters at runtime [73], while static quantization precomputes these parameters using calibration data [51]. Weight-only quantization applies only to model parameters [13], while full quantization includes both weights and activations [32].
The discrete nature of quantized representations introduces step-wise perturbations that may fundamentally alter the continuous structure of activation manifolds, making it particularly important to understand their topological implications [74].
@@
We implement post-training quantization [27, 37], mapping full-precision weights and activations to lower bit-widths using uniform quantization. We examine 8-bit, 4-bit, and 2-bit quantization schemes and their impact on the stability of topological features. Specifically, we evaluate INT8 Dynamic Quantization. It reduces the precision of linear layer weights to 8-bit integers during inference while keeping activations in full precision. This represents a practical deployment scenario that balances compression efficiency with accuracy preservation. Our analysis focuses on how quantization-induced discretization affects the continuity of activation spaces and the stability of persistent homological features. We particularly investigate whether the step-wise nature of quantized representations creates artificial clustering patterns or disrupts meaningful topological structures learned during training; in practice, quantization tends to snap nearby points (altering local neighborhoods), suppress low-variance directions, and sever thin connectors, which shifts several $H_{0}$ birth–death orderings. In Fig. 9, the “after” view shows widespread cluster breakup, spurious overlaps, and the disappearance of bridging structure rather than benign tightening.
@@ 6 Future Work
Current distance formulations assume Euclidean geometry in activation space. We propose investigating geodesic distances on learned manifolds and developing density-adaptive metrics that account for local neighborhood structure in high-dimensional embeddings [46]. Systematic investigation of topological changes under adversarial perturbations [68] may reveal fundamental principles of adversarial vulnerability and inform the development of topology-based defense mechanisms.
@@
Recurrent and transformer architectures process sequential information through complex temporal dependencies. We propose developing formulations that can capture the evolution of topological features across sequence positions, providing insights into memory formation and temporal representation learning. Extending our framework to generative architectures presents unique opportunities. Analysis of latent space topology in variational autoencoders [36] and the generator manifold in generative adversarial networks [25] could provide insights into mode collapse, sample diversity, and generation quality.
@@
Adapting methods from spatial data analysis [64], we propose incorporating kernel density estimation for persistence diagram visualization. This approach would identify regions of high topological activity and provide intuitive representations of feature density distributions. Computational complexity represents a significant barrier to applying topological analysis to large networks and datasets. We will consider investigating coreset approximations [20] that maintain topological fidelity while reducing computational requirements, enabling analysis of production-scale models.

### v2

@@ 4.5.1 Pruning
Pruning is a model compression technique that helps sparsify a network by systematically removing redundant or less important parameters by setting them to zero [lecun1989optimal, hassibi1993second]. The fundamental principle behind pruning is based on the observation that many neural networks are over-parameterized [frankle2018lottery], containing numerous weights that contribute minimally to the final output. By identifying and removing these parameters, we can achieve significant reductions in model size and computational requirements while maintaining comparable performance.
Pruning typically involves three main steps. First, during Importance Scoring, each parameter is assigned a score reflecting its contribution to the network’s performance. Next, during Selection, parameters with scores below a certain threshold are marked for removal. Finally, during Removal, the selected parameters are set to zero or completely eliminated from the network structure [blalock2020state].
There are several variants of pruning strategies. With magnitude-based pruning, parameters are ranked by their absolute values, with smaller weights considered less important and removed first [han2015deep]. Structured pruning removes entire channels, filters, or layers [li2016pruning, liu2017learning], while unstructured pruning removes individual weights regardless of their position in the network architecture [louizos2017learning]. Finally, global pruning considers all parameters across the entire network when making removal decisions [han2015deep], while layer-wise pruning makes decisions independently within each layer [molchanov2016pruning].
@@
We apply magnitude-based unstructured pruning [han2015deep] to remove weights below a threshold $\tau$, effectively setting $w_{ij}=0$ for $|w_{ij}|<\tau$. This sparsification process reduces the effective dimensionality of activation spaces and may eliminate critical topological connections. We evaluate pruning ratios from 10% to 90% and analyze how the resulting sparse networks maintain or lose persistent homological features across layers. We implement two distinct pruning strategies to analyze their differential impact on topological structure:
@@ 4.5.2 Quantization
Quantization is a compression technique that reduces the precision of numerical representations used in neural networks, typically converting floating-point weights and activations to lower bit-width representations [gholami2021survey]. This process significantly reduces memory requirements and computational complexity, making models more suitable for deployment on resource-constrained devices such as mobile phones and embedded systems [jacob2018quantization].
The quantization process involves mapping continuous floating-point values to a discrete set of quantized values [krishnamoorthi2018quantizing]. The most common approach is uniform quantization, where the range of possible values is divided into equally-spaced intervals. For a given floating-point value $x$, the quantized representation $q$ is computed as: $q=\text{round}\left(\frac{x-z}{s}\right)$, where $s$ is the scale factor that determines the spacing between quantized levels, and $z$ is the zero-point that ensures the quantization scheme can represent zero exactly [jacob2018quantization].
Several quantization strategies exist. Post-training quantization is applied after the model has been fully trained, requiring no retraining but potentially leading to accuracy degradation [banner2019post, nagel2019data]. Quantization-aware training incorporates quantization effects during the training process, typically achieving better accuracy preservation [jacob2018quantization, esser2019learned]. Dynamic quantization determines scale and zero-point parameters at runtime [wu2020integer], while static quantization precomputes these parameters using calibration data [migacz2017nvidia]. Weight-only quantization applies only to model parameters [dettmers2022gpt3], while full quantization includes both weights and activations [jacob2018quantization].
The discrete nature of quantized representations introduces step-wise perturbations that may fundamentally alter the continuous structure of activation manifolds, making it particularly important to understand their topological implications [yao2021hawq].
@@
We implement post-training quantization [han2015deep, krishnamoorthi2018quantizing], mapping full-precision weights and activations to lower bit-widths using uniform quantization. We examine 8-bit, 4-bit, and 2-bit quantization schemes and their impact on the stability of topological features. Specifically, we evaluate INT8 Dynamic Quantization. It reduces the precision of linear layer weights to 8-bit integers during inference while keeping activations in full precision. This represents a practical deployment scenario that balances compression efficiency with accuracy preservation. Our analysis focuses on how quantization-induced discretization affects the continuity of activation spaces and the stability of persistent homological features. We particularly investigate whether the step-wise nature of quantized representations creates artificial clustering patterns or disrupts meaningful topological structures learned during training; in practice, quantization tends to snap nearby points (altering local neighborhoods), suppress low-variance directions, and sever thin connectors, which shifts several $H_{0}$ birth–death orderings. In Fig. 9, the “after” view shows widespread cluster breakup, spurious overlaps, and the disappearance of bridging structure rather than benign tightening.
@@ 6 Future Work
Current distance formulations assume Euclidean geometry in activation space. We propose investigating geodesic distances on learned manifolds and developing density-adaptive metrics that account for local neighborhood structure in high-dimensional embeddings [lou2020neural]. Systematic investigation of topological changes under adversarial perturbations [szegedy2013intriguing] may reveal fundamental principles of adversarial vulnerability and inform the development of topology-based defense mechanisms.
@@
Recurrent and transformer architectures process sequential information through complex temporal dependencies. We propose developing formulations that can capture the evolution of topological features across sequence positions, providing insights into memory formation and temporal representation learning. Extending our framework to generative architectures presents unique opportunities. Analysis of latent space topology in variational autoencoders [kingma2013auto] and the generator manifold in generative adversarial networks [goodfellow2014generative] could provide insights into mode collapse, sample diversity, and generation quality.
@@
Adapting methods from spatial data analysis [silverman2018density], we propose incorporating kernel density estimation for persistence diagram visualization. This approach would identify regions of high topological activity and provide intuitive representations of feature density distributions. Computational complexity represents a significant barrier to applying topological analysis to large networks and datasets. We will consider investigating coreset approximations [feldman2020turning] that maintain topological fidelity while reducing computational requirements, enabling analysis of production-scale models.
Acknowledgements.
The authors wish to thank A, B, and C. This work was supported in part by a grant from XYZ.